What is the slope of the tangent line to the curve y = 2x^4 + x^3 + 6x at x = 1
Answer by Alan3354(69443)

What is the slope of the tangent line to the curve y = 2x^4 + x^3 + 6x at x = 1
---------
y' = 8x^3 + 3x^2 + 6
At x = 1, y' = 8 + 3 + 6 = 17
Slope = 17